OpenStudy (anonymous):

How do you turn 2x-3y=6 into y=mx+b form?

OpenStudy (anonymous):

@ChantelS any ideas?

OpenStudy (anonymous):

2x-3y=6 Subtract 2 on both sides. -3y = -2x+6 Divide both sides by -3. y = 2/3x+2

OpenStudy (anonymous):

sorry i meant y = 2/3x -2

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

Have a great day :D

OpenStudy (anonymous):

you too